Costs associated with the manufacture of miniature high-sensitivity piezoresistive pressure transducers is $74,000 per year. A clever industrial engineer found that by spending $17,000 now to reconfigure the production line and reprogram two of the robotic arms, the cost will go down to $56,000 next year and $51,000 in years 2 through 5. Using an interest rate of 14% per year, determine the present worth of the savings due to the reconfiguration.
Answer #1

Cost of reconfiguration =$17000

Savings in year 1 = 74000 - 56000 = $18000

Saving in year 2 to 5 = 74000 - 51000 = $23000

R = 14%

So,

PW of savings = 18000/1.14 + (23000*(1-1/1.14^4)/.14)*(1/1.14) - 17000

PW of savings = $57574.9 or $57575

So, the present worth of savings is $57574.9 or $57575.